OpenStudy (anonymous):

for every 3 boys standing in a line, there are 4 girls

OpenStudy (anonymous):

so whats your question?

OpenStudy (anonymous):

\[\frac{3}{3+4}=\frac{3}{7}=43\% \text{ are} \text{ boys}\]Is the above the question?